• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Epstein

    Log entry time 00:14:59 on March12,2000

    Entry number 36163

    keyword=spectrometer fields reset

    Resetting the spectrometers: the hadron arm seems to be coming back almost
    exactly as it was before. The electron arm dipole was set using probe 0 with
    regulation on. This brought the momentum setting as shown on the stip chart
    to within 20 KeV of where it was before the ESR shutdown. The current in electron Q1 was adjusted to come back to its old value which also brought it to the desired momentum setting of 0.417 GeV/c. (Note we had been using probe 0 for the electron dipole for almost all of the Eo=1.254 data taking. When the electron dipole was first set probe 3 was used, but a few hours later John LeRose changed to probe 0 and it stayed there from that point on.)